# KeyNode-Driven Geometry Coding for Real-World Scanned Human Dynamic Mesh Compression

### v1

@@ Abstract
The compression of real-world scanned 3D human dynamic meshes is an emerging research area, driven by applications such as telepresence, virtual reality, and 3D digital streaming. Unlike synthesized dynamic meshes with fixed topology, scanned dynamic meshes often not only have varying topology across frames but also scan defects such as holes and outliers, increasing the complexity of prediction and compression. Additionally, human meshes often combine rigid and non-rigid motions, making accurate prediction and encoding significantly more difficult compared to objects that exhibit purely rigid motion. To address these challenges, we propose a compression method designed for real-world scanned human dynamic meshes, leveraging embedded key nodes. The temporal motion of each vertex is formulated as a distance-weighted combination of transformations from neighboring key nodes, requiring the transmission of solely the key nodes’ transformations. To enhance the quality of the KeyNode-driven prediction, we introduce an octree-based residual coding scheme and a Dual-direction prediction mode, which uses I-frames from both directions. Extensive experiments demonstrate that our method achieves significant improvements over the state-of-the-art, with an average bitrate saving of 24.51% across the evaluated sequences, particularly excelling at low bitrates.
@@ I Introduction
  • Tackling Topology Variability: Instead of calculating the motion vector of individual vertices, which can be highly sensitive to changes in topology and scan defects, our method uses the motion of sparse controlling keynodes to represent vertex movement. This approach effectively captures the dynamics of the mesh, even with varying topology.
@@
  • Effective Temporal Prediction Enhancement: We propose Octree-based residual coding and a Dual-Direction prediction mode to effectively address prediction errors and better capture the complexities of dynamic human motion over time.
@@ II Related Work
Video-based methods: The Visual Volumetric Video-based Coding (V3C) standard is used in [17] to encode meshes with orthogonal projections, followed by atlas packing and video coding. Numerous proposals employing video-based methods were put forward in response to MPEG’s CfP for Dynamic Mesh Compression [18]. MPEG’s benchmark codec, known as Anchor, involves initial geometry encoding of each frame with mesh decimation and Draco compression [19]. Texture frames are stacked into a video and encoded with standard video coding techniques. A patch-based approach was presented in [20] reorganizing the intra-frame and inter-frame texture tiles to improve spatial and temporal correlation. In addition to mesh decimation, Mammou et al. proposed incorporating mesh subdivision, producing a base mesh and a series of subdivided meshes linked to their respective displacement vectors [21]. The displacements undergo wavelet transform, followed by quantization and packing into a 2D image or video, which is encoded by an image/video compression technique. MPEG adopted this approach [21] for further development into a new standard, named Video-based Dynamic Mesh Coding (V-DMC). Following its establishment, various efforts have been made to enhance V-DMC [22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
@@ III Proposed Codec
To address the challenge of inconsistent topology, which eliminates the need for explicit vertex correspondence, our KeyNode-Driven Compression method, originally introduced in [1], leverages embedded key nodes to represent motion vectors while ensuring efficient compression. The temporal motion of each vertex is formulated as a distance-weighted combination of transformations from its neighboring key nodes. The process is depicted in Fig. 1. Starting with the geometry of frame $t-1$, we use the embedded key nodes to predict the geometry of frame $t$ through embedded deformation, where the geometry of frame $t-1$ is deformed to estimate the geometry of frame $t$. Since keynodes are typically sparse and the motion vectors are quantized, prediction errors arise. To address these, we employ Octree-based residual coding to refine the predicted geometry for frame $t$. At the sequence level, we enhance the prediction process with a Dual-direction prediction mode. The details of the KeyNode-Driven coding are presented in Section III-A, residual coding in Section III-B, and the Dual-direction prediction in Section III-C.
@@ III-A KeyNode-Driven Geometry Coding
Consider a dynamic mesh $S=\{M_{1},M_{2},...,M_{F}\}$, where $F$ represents the total number of frames in the sequence, and $M_{t}$ denotes a static mesh at time $t$ containing $V_{t}$ vertices. The average number of vertices is denoted $V$. Static meshes in $S$ may exhibit varying topologies.
Given two mesh frames $M_{t-1}$ and $M_{t}$, an approximation of $M_{t}$, denoted $M^{\prime}_{t}$, can be obtained through deforming $M_{t-1}$ using an appropriate set of key nodes and their transformations. Let $\mathcal{GT}(.)$ be a geometric transformation operator which deforms $M_{t-1}$ to generate $M^{\prime}_{t}$:
@@
Here, $\textbf{n}=\{n_{j}\in\mathbb{R}^{3}\}\in\mathbb{R}^{N\times 3}$ is the set of key nodes, $\textbf{R}=\{R_{j}\in\mathbb{R}^{3}\}\in\mathbb{R}^{N\times 3}$ and $\textbf{T}=\{t_{j}\in\mathbb{R}^{3}\}\in\mathbb{R}^{N\times 3}$ are the key nodes’ affine transformations (rotations and translations), with $j\in[1,N]$ where $N<\!\!<V$ is the total number of key nodes controlling the deformation of the mesh.
@@
Key node placement impacts deformation quality. In [1], we presented an efficient optimization solution that balances compactness and quality by identifying the optimal set of key nodes. We formulated the balancing goal as a dual-objective optimization problem using the Alternating Direction Method of Multipliers (ADMM). To simplify the dual-objective formulation, a structured three-step approach (comprising a mesh registration model, coordinate descent, and gradient descent) is iteratively applied in a loop until the desired node sparsity is achieved. The Optimal Key Node Generator block in Fig. 2 employs that method, ensuring the set of chosen key nodes meets the target sparsity while minimizing distortion.
@@
Here, $\mathcal{L}_{data}$ penalizes the geometry deviation between source and target; $\mathcal{L}_{rot}$ measures the deviation of Rotation R from the orthogonal matrix; and $\mathcal{L}_{reg}$ ensures the smoothness of the deformation. We employ the approach introduced in [32] to solve this optimization problem.
For each P-frame, keynode transformations including $\textbf{R}\in\mathbb{R}^{N\times 3}$ and $\textbf{T}\in\mathbb{R}^{N\times 3}$ are flattened into a rotation vector and translation vector, denoted $\textbf{R}^{f}\in\mathbb{R}^{3N}$ and $\textbf{T}^{f}\in\mathbb{R}^{3N}$. Two Cauchy distributions are fitted to $\textbf{R}^{f}$ and $\textbf{T}^{f}$. With a specified codebook size $N^{RT}_{b}$, $N^{RT}_{b}$ equal-width bins are generated spanning from $-|b|$ to $|b|$, where $b$ is the maximum absolute value within a given frame. The two middle bins are merged to create a dead-zone. The bin $(b_{1}^{k},b_{2}^{k}]$ has its quantized value as the expected value of the Cauchy probability density function (pdf) within that range:
@@ III-B Octree-based Geometric Residual Coding
Predicting P-frames using embedded key nodes introduces distortion, so we use a Geometric Residual Coding method based on octrees to reduce this geometric distortion.
@@ III-B2 Cost-constrained Octree Generation
Determining a suitable octree structure to efficiently quantize the residuals is crucial. In this section, we present a cost-constrained approach for generating a suitable unbalanced octree for each residual frame, balancing residual distortion and octree size (bitrate) through a cost function of an optimization problem. When a residual frame $\mathbf{r_{t}}$ is quantized by an octree $T$, we denote the quantized frame by $\mathbf{\hat{r}_{t}}=\{\hat{r}_{0},\hat{r}_{1},...,\hat{r}_{V_{IF}-1}\}$, where each $\hat{r}_{i}$ represents the average residual of the leaf node to which $r_{i}$ belongs. The overall quantization distortion can be calculated as the mean squared error between the original and quantized residual frames:
@@
Here, $D(T)$ is the distortion of the octree $T$ in which $n$ is a leaf node, $D(T(n))$ is the distortion of the octree in which $n$ is expanded, and $\lambda(n)$ is the slope of the cost-distortion curve when adjusting the octree structure.
When growing the octree, the node with the largest $\lambda(n)$ will be chosen to expand, and when pruning, the node with the smallest $\lambda(n)$ will be made a leaf.
To avoid calculating the distortion and cost for the whole tree, we compute $\lambda(n)$ based on the distortion of the node being evaluated [33]:
@@ III-B3 Coding of Octree-based Quantized Residuals
Fig. 6 depicts the structure for residual coding. At the encoder, the distorted P-frame $M^{\prime}_{t}$ is used to generate a balanced octree based on the vertex positions. When vertices are accurately predicted through deformation, they should not be corrected with residual coding. To determine which nodes should carry residuals, we compare the average vertex distortion $e^{l}$ of the node’s vertices with a threshold $\mathcal{T}$. In our implementation, the value of $e^{l}$ is derived from the Mesh Structural Distortion Measure ver. 2 (MSDM2) [34] distortion metric, while it could also be directly calculated as $\hat{r}^{l}$ to save computational cost. All nodes where $e^{l}<\mathcal{T}$ are referred to as Non-Correcting Octree Cells (NCOC); residuals for their vertices will not be transmitted. Non-correcting flags are sent only for the highest-level NCOC parents, as their children inherit the same type. The octree is then pruned using the cost-constrained method. Pruning flags are sent to the decoder, allowing it to reconstruct the unbalanced octree.
@@
The residuals are quantized with the unbalanced octree, and then each dimension of $\textbf{r}^{\mathcal{L}}$ undergoes uniform quantization with a dead-zone. A Cauchy distribution is fitted to the original residual frame $\mathbf{r_{t}}$ in the corresponding dimension, which is used to construct the Huffman codebook in a manner similar to the process described in Eqs. 4 and 5. This process is repeated for each dimension independently. Overall, transmitting the residual frame $\mathbf{r_{t}}$ requires: two binary streams of flags, three bitstreams of Huffman-coded leaf residuals, and three fitted Cauchy distributions described by their means and variances.

### v2

@@ Abstract
The compression of real-world scanned 3D human dynamic meshes is an emerging research area, driven by applications such as telepresence, virtual reality, and 3D digital streaming. Unlike synthesized dynamic meshes with fixed topology, scanned dynamic meshes often not only have varying topology across frames but also scan defects such as holes and outliers, increasing the complexity of prediction and compression. Additionally, human meshes often combine rigid and non-rigid motions, making accurate prediction and encoding significantly more difficult compared to objects that exhibit purely rigid motion. To address these challenges, we propose a compression method designed for real-world scanned human dynamic meshes, leveraging embedded key nodes. The temporal motion of each vertex is formulated as a distance-weighted combination of transformations from neighboring key nodes, requiring the transmission of solely the key nodes’ transformations. To enhance the quality of the KeyNode-driven prediction, we introduce an octree-based residual coding scheme and a Dual-direction prediction mode, which uses I-frames from both directions. Extensive experiments demonstrate that our method achieves significant improvements over the state-of-the-art, with an average bitrate savings of 58.43% across the evaluated sequences, particularly excelling at low bitrates.
@@ I Introduction
  • A Novel Approach to Vertex Motion Modeling: We propose a KeyNode-driven framework for modeling vertex motion, where each vertex is influenced by a weighted combination of transformations from a sparse set of key nodes. Unlike block-based methods, which assign a single rigid motion vector to entire spatial regions, our approach allows fine-grained, non-rigid deformation with per-vertex flexibility. This enables more accurate modeling of complex motion while maintaining a compact representation.
  • Tackling Topology Variability: Instead of calculating the motion vector of individual vertices, which can be highly sensitive to changes in topology and scan defects as well as costly in bitrate, our method uses the motion of sparse controlling keynodes to represent vertex movement. This approach effectively captures the dynamics of the mesh, even with varying topology.
@@
  • Compression-aware Key Node Optimization: Our work is the first to introduce a framework for selecting both the number and position of embedded key nodes in order to balance geometric fidelity with transmission cost. Prior embedded deformation work focused only on deformation quality, with no consideration of bitrate.
  • Residual-aware Octree Design: Our KeyNode-driven prediction framework produces residuals that naturally cluster in localized surface regions. To exploit this structure, we propose an optimization scheme for constructing an unbalanced octree tailored to the spatial distribution of geometric residuals. Our approach dynamically balances bitrate and reconstruction accuracy by selectively expanding or pruning tree nodes.
@@ II Related Work
Video-based methods: The Visual Volumetric Video-based Coding (V3C) standard is used in [17] to encode meshes with orthogonal projections, followed by atlas packing and video coding. Numerous proposals employing video-based methods were put forward in response to MPEG’s CfP for Dynamic Mesh Compression [18]. MPEG’s benchmark codec, known as Anchor, involves initial geometry encoding of each frame with mesh decimation and Draco compression [19]. Texture frames are stacked into a video and encoded with standard video coding techniques. A patch-based approach was presented in [20] reorganizing the intra-frame and inter-frame texture tiles to improve spatial and temporal correlation. In addition to mesh decimation, Mammou et al. proposed incorporating mesh subdivision, producing a base mesh and a series of subdivided meshes linked to their respective displacement vectors, and the temporal motion is captured by tracking vertex-level positional differences between consecutive base meshes [21]. The displacements undergo wavelet transform, followed by quantization and packing into a 2D image or video, which is encoded by an image/video compression technique. MPEG adopted this approach [21] for further development into a new standard, named Video-based Dynamic Mesh Coding (V-DMC). Following its establishment, various efforts have been made to enhance V-DMC [22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
@@ III Proposed Codec
As outlined in the Related Work section, many existing methods model temporal motion in dynamic mesh sequences by calculating per-block positional differences between successive frames. This strategy is conceptually similar to motion estimation in 2D video compression standards in H.264 or MPEG, where motion is often modeled using block-based techniques, assuming uniform, unidirectional motion across a region, with a single motion vector shared by all pixels within a block. Other existing methods get around this limitation by calculating per-vertex positional differences. However, this method would significantly increase the bit rate for transmitting vertex-level motion vectors.
Our approach, by contrast, adopts Embedded Deformation (ED) to model vertex-level motion in 3D space. ED enables smooth, non-rigid, and multi-directional motion by expressing each vertex’s movement as a weighted combination of transformations from a sparse set of key nodes. Because a small number of key nodes can efficiently encode complex deformations, this sparse key node approach can avoid the rate penalty of per-vertex motion information, but it provides more flexibility than the per-block methods, as each vertex can move independently without being bound to rigid spatial groupings, as in the case for block-based models. Importantly, this approach tackles core challenges specific to scanned dynamic 3D meshes, particularly the absence of reliable vertex-to-vertex correspondence between frames in real-world 3D scans. Unlike 2D pixels, which maintain fixed grid positions and indices, vertices in dynamic meshes often vary in topology and ordering over time. This makes simple frame-to-frame differencing unreliable. Our deformation-based model overcomes these limitations, offering a more robust and generalizable solution for capturing temporal motion in dynamic 3D geometry.
Following our preliminary work in [1], which introduced the use of ED to model vertex-level motion in dynamic meshes, a recent paper [23] also proposed using ED for motion modeling. While both works adopt ED as a core concept, their approach differs from ours in two important ways. First, their usage of ED is limited to predicting the motion of a simplified base mesh generated from decimation, which is later subdivided to recover geometric detail. In contrast, we leverage the full flexibility of ED to model motion directly at the vertex level, capturing fine-grained deformations without relying on post-subdivision. Second, as noted in their paper, their method is intended for use as a preprocessing step outside the core V-DMC codec. As such, it does not incorporate any bitrate constraints or optimize for compression. In contrast, our approach is explicitly designed for integration into the compression pipeline, so the design of the embedded deformation itself directly balances the competing goals of motion fidelity and transmission cost.
The process is depicted in Fig. 1. Starting with the geometry of frame $t-1$, we use the embedded key nodes to predict the geometry of frame $t$ through embedded deformation, where the geometry of frame $t-1$ is deformed to estimate the geometry of frame $t$. Since keynodes are typically sparse and the motion vectors are quantized, prediction errors arise. To address these, we employ Octree-based residual coding to refine the predicted geometry for frame $t$. At the sequence level, we enhance the prediction process with a Dual-direction prediction mode. The details of the KeyNode-Driven coding are presented in Section III-A, residual coding in Section III-B, and the Dual-direction prediction in Section III-C.
@@ III-A KeyNode-Driven Geometry Coding
Consider a dynamic mesh $S=\{M_{1},M_{2},...,M_{F}\}$, where $F$ represents the total number of frames in the sequence, and $M_{t}$ denotes a static mesh at time $t$ containing $V_{t}$ vertices. The average number of vertices is denoted $V$. Static meshes in $S$ may exhibit varying topologies. Given two mesh frames $M_{t-1}$ and $M_{t}$, an approximation of $M_{t}$, denoted $M^{\prime}_{t}$, can be obtained through deforming $M_{t-1}$ using an appropriate set of key nodes and their transformations. Let $\mathcal{GT}(.)$ be a geometric transformation operator which deforms $M_{t-1}$ to generate $M^{\prime}_{t}$:
@@
Here, $\textbf{n}=\{n_{j}\in\mathbb{R}^{3}\}\in\mathbb{R}^{N\times 3}$ is the set of key nodes, $\textbf{R}=\{R_{j}\in\mathbb{R}^{3}\}\in\mathbb{R}^{N\times 3}$ and $\textbf{T}=\{t_{j}\in\mathbb{R}^{3}\}\in\mathbb{R}^{N\times 3}$ are the key nodes’ transformations (rotations and translations), with $j\in[1,N]$ where $N<\!\!<V$ is the total number of key nodes controlling the deformation of the mesh.
@@
The placement of key nodes plays a critical role in the quality of deformation. While ED has been widely adopted in geometry processing and animation tasks, previous works typically assume a fixed, predefined set of key nodes and focus solely on improving deformation accuracy. The key nodes are determined by uniform sampling over the mesh surface in [10.1145/1276377.1276478] and [32], while [23] uses the base mesh, a decimated version of the original mesh. In the context of compression, however, it is essential to also consider bitrate, introducing a new trade-off between geometric fidelity and representation cost.
To tackle this, our earlier work [1] introduced a Key Node Generator, which formulates an optimization framework to jointly balance deformation quality and transmission efficiency. A central challenge here is that incorporating bitrate, which ties closely to the number of key nodes, directly into the standard ED optimization is problematic, as ED traditionally assumes the key nodes are fixed when solving for the local transformations (R and T). Our method circumvents this by alternating between estimating deformations, pruning key nodes, and adjusting key node positions. This allows the layout of key nodes to adapt dynamically based on both the underlying motion and the associated encoding cost. To the best of our knowledge, this is the first approach to integrate ED with a compression-aware optimization of the key node configuration, enabling a more efficient representation of dynamic 3D geometry. The Optimal Key Node Generator block in Fig. 2 employs that method, ensuring the set of chosen key nodes meets the target sparsity while minimizing distortion.
@@
Here, $\mathcal{L}_{data}$ penalizes the geometry deviation between source and target, and $\mathcal{L}_{reg}$ ensures the smoothness of the deformation. We employ the approach introduced in [32] to solve this optimization problem. Note that the approach in [32] is designed to maximize deformation quality without considering compression constraints. To this end, it employs a densely sampled set of key nodes distributed uniformly over the mesh surface. In contrast, our work addresses the fundamentally different challenge of compression, where bitrate plays a major role. Rather than using a dense and fixed key node layout, we replace it with a sparse, optimized key node set generated by our previously introduced Optimal Key Node Generator. This compression-aware design balances accuracy with representation cost, enabling efficient and scalable mesh prediction.
For each P-frame, keynode transformations including $\textbf{R}\in\mathbb{R}^{N\times 3}$ and $\textbf{T}\in\mathbb{R}^{N\times 3}$ are flattened into a rotation vector and translation vector, denoted $\textbf{R}^{f}\in\mathbb{R}^{3N}$ and $\textbf{T}^{f}\in\mathbb{R}^{3N}$. Fig. 3 presents the distributions of our data with the fitted Gaussian, Laplace, and Cauchy distributions. Among these, the Cauchy distribution demonstrates the closest alignment with the observed data, capturing both the sharp peak and heavy tails more accurately than the others. This insight plays a key role in our compression strategy. By modeling our transformations with a Cauchy distribution, we are able to design a quantization scheme that is statistically well-matched to the data. Crucially, this also allows for the implicit construction of the Huffman coding table at the decoder—meaning the coding dictionary does not need to be transmitted explicitly. This eliminates additional overhead and enhances compression efficiency.
Figure 3: Distributions of RT and Residuals for frame 619 of Thomas. The histograms are plotted with 200 bins.
Two Cauchy distributions are fitted to $\textbf{R}^{f}$ and $\textbf{T}^{f}$. With a specified codebook size $N^{RT}_{b}$, $N^{RT}_{b}$ equal-width bins are generated spanning from $-|b|$ to $|b|$, where $b$ is the maximum absolute value within a given frame. The two middle bins are merged to create a dead-zone. The bin $(b_{1}^{k},b_{2}^{k}]$ has its quantized value as the expected value of the Cauchy probability density function (pdf) within that range:
@@ III-B Octree-based Geometric Residual Coding
Residual coding is a critical component of any predictive compression method. Unlike prior mesh compression approaches such as V-DMC—which flatten 3D vertex-level residuals into 2D images and encode them using conventional video compression techniques—our KeyNode-based prediction framework operates natively in 3D space. We observed that the geometric residuals naturally cluster in localized surface regions, due to the spatial smoothness and locality of the deformation model.
This spatial coherence of residuals motivates the proposed octree-structured residual coding. The octree structure allows residuals to be encoded at varying spatial resolutions, adapting to the complexity of different regions. Importantly, we do not use a standard balanced octree. Instead, we propose an optimization approach to determine an optimal unbalanced octree, which considers both correction quality and bitrate during construction.
@@ III-B2 Cost-constrained Octree Generation
Determining a suitable octree structure to efficiently quantize the residuals is crucial. In this section, we present a cost-constrained optimization approach for generating a suitable unbalanced octree for each residual frame, balancing residual distortion and octree size (bitrate) through a cost function of an optimization problem. When a residual frame $\mathbf{r_{t}}$ is quantized by an octree $T$, we denote the quantized frame by $\mathbf{\hat{r}_{t}}=\{\hat{r}_{0},\hat{r}_{1},...,\hat{r}_{V_{IF}-1}\}$, where each $\hat{r}_{i}$ represents the average residual of the leaf node to which $r_{i}$ belongs. The overall quantization distortion can be calculated as the mean squared error between the original and quantized residual frames:
@@
Here, $D(T)$ is the distortion of the octree $T$ in which $n$ is a leaf node, $D(T(n))$ is the distortion of the octree in which $n$ is expanded, and $\lambda(n)$ is the slope of the cost-distortion curve when adjusting the octree structure. When growing the octree, the node with the largest $\lambda(n)$ will be chosen to expand, and when pruning, the node with the smallest $\lambda(n)$ will be made a leaf. To avoid calculating the distortion and cost for the whole tree, we compute $\lambda(n)$ based on the distortion of the node being evaluated [33]:
@@ III-B3 Coding of Octree-based Quantized Residuals
Fig. 7 depicts the structure for residual coding. At the encoder, the distorted P-frame $M^{\prime}_{t}$ is used to generate a balanced octree based on the vertex positions. When vertices are accurately predicted through deformation, they should not be corrected with residual coding. To determine which nodes should carry residuals, we compare the average vertex distortion $e^{l}$ of the node’s vertices with a threshold $\mathcal{T}$. In our implementation, the value of $e^{l}$ is derived from the Mesh Structural Distortion Measure ver. 2 (MSDM2) [34] distortion metric, while it could also be directly calculated as $\hat{r}^{l}$ to save computational cost. All nodes where $e^{l}<\mathcal{T}$ are referred to as Non-Correcting Octree Cells (NCOC); residuals for their vertices will not be transmitted. An octree cell is also classified as an NCOC if its local MSE distortion after residual correction is greater than or equal to the MSE distortion before correction. Non-correcting flags are sent only for the highest-level NCOC parents, as their children inherit the same type. The octree is then pruned using the cost-constrained optimization method. Pruning flags are sent to the decoder, allowing it to reconstruct the unbalanced octree. We also disable residual coding for the entire frame if the distortion increases after correction.
@@
The residuals are quantized with the unbalanced octree, and then they undergo Cauchy-fitted uniform quantization with a dead-zone, similar to the way the R and T transformations are handled. A Cauchy distribution is fitted to the original residual frame $\mathbf{r_{t}}$ in the corresponding dimension, which is used to construct the Huffman codebook in a manner similar to the process described in Eqs. 4 and 5. Overall, transmitting the residual frame $\mathbf{r_{t}}$ requires: two binary streams of flags, a bitstream of Huffman-coded leaf residuals, and a fitted Cauchy distribution described by its mean and variance.
@@
Although our residual estimation process is similar to that of V-DMC—both compute residuals by projecting onto the original mesh surface—our method fundamentally diverges in how residuals are quantized and encoded. V-DMC applies a wavelet transform to the residuals (referred to as displacements), quantizes the coefficients, and encodes them by traditional video compression techniques after packing them into a 2D video. In contrast, our method maintains the intrinsic 3D nature of the data throughout the compression pipeline. We perform quantization directly in 3D space, leveraging an adaptive octree structure that aligns with the spatial distribution and complexity of the residuals. For entropy coding, we exploit the nature of the residuals’ distribution using a Cauchy-based coding scheme, enabling compact and statistically informed representation. By staying within the 3D domain, our method preserves geometric fidelity and achieves efficient compression.